I'm trying to find the x intercept of y=1/2x+1. Need help please
Answer by stanbon(75887)

I'm trying to find the x intercept of y=1/2x+1. Need help please
---
The x-intercept is the y value when x = 0.
-----
y = (1/2)x + 1
----
0 = (1/2)x + 1
-----
(1/2)x = -1
-----
x = -2
------
Ans: x-intercept is (-2,0)
------------------------------
Cheers,
Stan H.
--------------